OpenStudy (anonymous):

colored copies cost .10 cents white copies cost .5cents sandy make 400 copies , some of which were colored and the rest white. her total was $22.50. how many colored copies and how many white did she made?

myininaya (myininaya):

let colored copies be represented by C let white copies be represented by W .10C+.5W=22.50 C+W=400

myininaya (myininaya):

sine C+W=400, then W=400-C. so .10C+.5(400-C)=22.50 .10C+200-.5C=22.50 -.4C+200=22.50 -.4C=22.50-200=-177.5 C=-177.5/(-.4)=443.75 But thats impossible since Sandy only made 400 copies

myininaya (myininaya):

did you mean white copies cost .5 cents or .05 cents?

OpenStudy (anonymous):

.05 cents

myininaya (myininaya):

ok so we have .10C+.05W=22.50 C+W=400 implies W=400-C so .10C+.05(400-C)=22.50 .10C+20-.05C=22.50 .05C+20=22.50 .05C=22.50-20=2.5 C=2.5/(.05)=50 since W=400-C, then W=400-50=350

OpenStudy (anonymous):

i did it like this..c+w=400 .10c+5w=22.50 multiply#1 equation by .10 .10c-10w=17.50 add#3 equation and #2 400c=17.50 c=228 w+228=400

OpenStudy (anonymous):

but it doesn't match with the price:(

myininaya (myininaya):

ok did you understand how I got W=350 and C=50?

OpenStudy (anonymous):

no how???

myininaya (myininaya):

C+W=400 I solved for W and got W=400-C then I plugged this into the other equation

OpenStudy (anonymous):

ok now i got it thank you you save me from failing algebra :)

myininaya (myininaya):

lol k

OpenStudy (anonymous):

i got another ???

OpenStudy (anonymous):

an airplane flew for 5 hours with a tail wind of 25km/h. The return flight against the same wind took 5.5 hours. find the speed of the plane in still air